# Dependency Parsing is More Parameter-Efficient with Normalization

### v1

@@ B.6 Full fine-tuning and sample efficiency
In this section, we present the results obtained when fine-tuning BERTbase, DeBERTabase, BERTlarge, and DeBERTalarge over $10k$ steps. In our previous settings, we only used a frozen BERTbase, whose last hidden states we fed as input to the tagger and parser. This evaluation allows us to test the effectiveness of our approach in a setting in which the number of learnable parameters is unconstrained. In this experiment, we use different learning rates for the base models ($\eta=1\times 10^{-4}$) and the large models ($\eta=3\times 10^{-5}$) and we apply gradient norm clipping with $\|\nabla\|_{max}=1.0$. In the case of the large models, we also use a cosine schedule with warm-up over 6% of the steps. We adopt these measures because during early trials we experienced sudden mid-run gradient explosions. Note that, in this case, we do not use any downstream BiLSTMs and only vary whether we use biaffine score normalization in the parser.
@@
For both BERTbase and DeBERTabase, the difference in convergence speed and performance is statistically significant on ADE, CoNLL04, and SciERC. The same is true for BERTlarge only on ADE. However, the effect is statistically significant with DeBERTalarge for all datasets. This indicates our score normalization approach can produce positive effects in terms of sample efficiency also when fine-tuning models with hundreds of millions of parameters.
Table 9: Performance for the fully fine-tuned BERT and DeBERTa models ($h_{\psi}=400$, $h_{out}=500$, $\phi=\checkmark$ and $\textbf{e}_{i}^{tag}=\checkmark$).
@@
BERTbase $1$ 0.748 ±$\pm 0.028$ 0.613 ±$\pm 0.019$ 0.414 ±$\pm 0.011$ 0.726 ±$\pm 0.006$ 0.321
$\frac{1}{\sqrt{d}}$ 0.731 ±$\pm 0.025$ 0.629 ±$\pm 0.022$ 0.412 ±$\pm 0.016$ 0.726 ±$\pm 0.006$ 0.321
BERTlarge $1$ 0.748 ±$\pm 0.016$ 0.700 ±$\pm 0.019$ 0.473 ±$\pm 0.011$ 0.750 ±$\pm 0.006$ 0.340
$\frac{1}{\sqrt{d}}$ 0.777 ±$\pm 0.011$ 0.697 ±$\pm 0.014$ 0.446 ±$\pm 0.025$ 0.748 ±$\pm 0.010$ 0.341
DeBERTabase $1$ 0.754 ±$\pm 0.013$ 0.674 ±$\pm 0.014$ 0.429 ±$\pm 0.015$ 0.751 ±$\pm 0.002$ 0.332
$\frac{1}{\sqrt{d}}$ 0.761 ±$\pm 0.021$ 0.700 ±$\pm 0.013$ 0.425 ±$\pm 0.019$ 0.751 ±$\pm 0.010$ 0.338
DeBERTalarge $1$ 0.786 ±$\pm 0.010$ 0.739 ±$\pm 0.016$ 0.478 ±$\pm 0.019$ 0.763 ±$\pm 0.006$ 0.352
$\frac{1}{\sqrt{d}}$ 0.794 ±$\pm 0.015$ 0.747 ±$\pm 0.013$ 0.476 ±$\pm 0.010$ 0.763 ±$\pm 0.007$ 0.353
@@
Appendix C Computational resources
[…]
@@ 3. Answer: [Yes]
  4. Justification: We believe our abstract and introduction faithfully report the claims made in the rest of the paper, i.e. we show theoretically and empirically that normalizing biaffine scores for dependency parsing in NLP provides higher parsing performance.
@@ 23. Answer: [Yes]
  24. Justification: The code is provided. The repository includes a README.md file which guides the reader through its usage. The data for ADE, CoNLL04, and SciERC can be downloaded the link provided in Section 4.1. As regards ERFGC, [36] provide access to the corpus upon request. UD 2.2 enEWT and SciDTB are openly available.
@@ 28. Answer: [Yes]
  29. Justification: Yes, we report information about the data we use in Section 4.1. Training and evaluation information is provided in Sections 4.2, 4.3 and 4.4.
@@ 35. Guidelines:
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
@@ 38. Answer: [Yes]
  39. Justification: Appendix C provides all relevant information with relation to the hardware used, its characteristics, and the time necessary to reproduce the experiments.
@@ 43. Answer: [Yes]
  44. Justification: We believe our work does NOT raise any ethics concerns with regard to any of the guidelines listed within the NeurIPS Code of Ethics. We have no human participants, we use publicly available data, and we believe our work does not have any broader potentially unintended societal impact.
@@ 60. Guidelines:
  63. Answer: [N/A]
  64. Justification: We do not release any new assets.

### v2

@@ D.6 Full fine-tuning and sample efficiency
In this section, we present the results obtained when fine-tuning BERTbase, DeBERTabase, BERTlarge, and DeBERTalarge over $10k$ steps. In our previous settings, we only used a frozen BERTbase, whose last hidden states we fed as input to the tagger and parser. This evaluation allows us to test the effectiveness of our approach in a setting in which the number of learnable parameters is unconstrained. In this experiment, we use different learning rates for the base models ($\eta=1\times 10^{-4}$) and the large models ($\eta=3\times 10^{-5}$) and we apply gradient norm clipping with $\|\nabla\|_{max}=1.0$. In the case of the large models, we also use a cosine schedule with warm-up over 6% of the steps. We adopt these measures because during early trials we experienced sudden mid-run gradient explosions. In this case, we do not use any downstream BiLSTMs and only vary whether we use biaffine score normalization in the parser.
@@
For both BERTbase and DeBERTabase, the difference in convergence speed and performance is statistically significant on ADE, CoNLL04, and SciERC. The same is true for BERTlarge only on ADE. However, the effect is statistically significant with DeBERTalarge for all datasets. This indicates our score normalization approach can produce positive effects in terms of sample efficiency also when fine-tuning models with hundreds of millions of trainable parameters.
Table 12: Results for the fully fine-tuned models ($h_{\psi}=400$, $h_{out}=500$, $\phi=\checkmark$, and $\textbf{e}_{i}^{tag}=\checkmark$).
@@
BERTbase 11 0.748 $\pm 0.028$ 0.613 $\pm 0.019$ 0.414 $\pm 0.011$ 0.726 $\pm 0.006$ 0.321
$\frac{1}{\sqrt{d}}$ 0.731 $\pm 0.025$ 0.629 $\pm 0.022$ 0.412 $\pm 0.016$ 0.726 $\pm 0.006$ 0.321
BERTlarge 11 0.748 $\pm 0.016$ 0.700 $\pm 0.019$ 0.473 $\pm 0.011$ 0.750 $\pm 0.006$ 0.340
$\frac{1}{\sqrt{d}}$ 0.777 $\pm 0.011$ 0.697 $\pm 0.014$ 0.446 $\pm 0.025$ 0.748 $\pm 0.010$ 0.341
DeBERTabase 11 0.754 $\pm 0.013$ 0.674 $\pm 0.014$ 0.429 $\pm 0.015$ 0.751 $\pm 0.002$ 0.332
$\frac{1}{\sqrt{d}}$ 0.761 $\pm 0.021$ 0.700 $\pm 0.013$ 0.425 $\pm 0.019$ 0.751 $\pm 0.010$ 0.338
DeBERTalarge 11 0.786 $\pm 0.010$ 0.739 $\pm 0.016$ 0.478 $\pm 0.019$ 0.763 $\pm 0.006$ 0.352
$\frac{1}{\sqrt{d}}$ 0.794 $\pm 0.015$ 0.747 $\pm 0.013$ 0.476 $\pm 0.010$ 0.763 $\pm 0.007$ 0.353
@@
Table 13: Micro-F1 (SemDP) and LAS (SynDP) on all tasks ($\phi$ = ✓, $\textbf{e}_{i}^{tag}=\checkmark$). Best in bold.
Metric $a$ NN ADE CoNLL04 SciERC ERFGC enEWT SciDTB
$(h_{\psi},d_{\text{MLP}})=$ (200, 100) (400, 300) (300, 300) (400, 300) (400, 500) (400, 500)
rels. 11 0 0.541 $\pm 0.021$ 0.399 $\pm 0.024$ 0.147 $\pm 0.049$ 0.548 $\pm 0.010$ 0.559 $\pm 0.005$ 0.729 $\pm 0.004$
1 0.657 $\pm 0.011$ 0.556 $\pm 0.021$ 0.282 $\pm 0.009$ 0.676 $\pm 0.010$ 0.771 $\pm 0.006$ 0.892 $\pm 0.002$
2 0.667 $\pm 0.011$ 0.573 $\pm 0.025$ 0.273 $\pm 0.010$ 0.694 $\pm 0.010$ 0.796 $\pm 0.006$ 0.910 $\pm 0.002$
3 0.662 $\pm 0.027$ 0.562 $\pm 0.021$ 0.299 $\pm 0.023$ 0.705 $\pm 0.011$ 0.804 $\pm 0.006$ 0.915 $\pm 0.002$
$\frac{1}{\sqrt{d}}$ 0 0.567 $\pm 0.014$ 0.438 $\pm 0.033$ 0.181 $\pm 0.027$ 0.612 $\pm 0.008$ 0.646 $\pm 0.002$ 0.796 $\pm 0.002$
1 0.668 $\pm 0.017$ 0.597 $\pm 0.015$ 0.299 $\pm 0.019$ 0.692 $\pm 0.009$ 0.789 $\pm 0.003$ 0.904 $\pm 0.002$
2 0.676 $\pm 0.019$ 0.596 $\pm 0.014$ 0.312 $\pm 0.011$ 0.699 $\pm 0.009$ 0.805 $\pm 0.003$ 0.916 $\pm 0.002$
3 0.686 $\pm 0.025$ 0.602 $\pm 0.017$ 0.320 $\pm 0.013$ 0.708 $\pm 0.008$ 0.807 $\pm 0.005$ 0.919 $\pm 0.001$
edges 11 0 0.536 $\pm 0.009$ 0.415 $\pm 0.020$ 0.173 $\pm 0.050$ 0.601 $\pm 0.013$ 0.589 $\pm 0.006$ 0.745 $\pm 0.005$
1 0.652 $\pm 0.009$ 0.566 $\pm 0.022$ 0.321 $\pm 0.009$ 0.744 $\pm 0.013$ 0.793 $\pm 0.006$ 0.901 $\pm 0.002$
2 0.657 $\pm 0.021$ 0.586 $\pm 0.017$ 0.322 $\pm 0.017$ 0.769 $\pm 0.014$ 0.819 $\pm 0.006$ 0.919 $\pm 0.002$
3 0.649 $\pm 0.027$ 0.578 $\pm 0.011$ 0.355 $\pm 0.028$ 0.782 $\pm 0.007$ 0.827 $\pm 0.006$ 0.924 $\pm 0.002$
$\frac{1}{\sqrt{d}}$ 0 0.549 $\pm 0.014$ 0.453 $\pm 0.032$ 0.203 $\pm 0.030$ 0.674 $\pm 0.007$ 0.682 $\pm 0.003$ 0.815 $\pm 0.001$
1 0.654 $\pm 0.020$ 0.598 $\pm 0.021$ 0.351 $\pm 0.019$ 0.762 $\pm 0.009$ 0.810 $\pm 0.003$ 0.913 $\pm 0.002$
2 0.660 $\pm 0.015$ 0.600 $\pm 0.008$ 0.367 $\pm 0.015$ 0.775 $\pm 0.008$ 0.827 $\pm 0.003$ 0.925 $\pm 0.002$
3 0.666 $\pm 0.028$ 0.610 $\pm 0.022$ 0.377 $\pm 0.021$ 0.786 $\pm 0.004$ 0.829 $\pm 0.004$ 0.928 $\pm 0.002$
tags 11 0 0.615 $\pm 0.115$ 0.285 $\pm 0.150$ 0.014 $\pm 0.015$ 0.662 $\pm 0.073$
1 0.665 $\pm 0.147$ 0.671 $\pm 0.018$ 0.049 $\pm 0.014$ 0.794 $\pm 0.057$
2 0.746 $\pm 0.011$ 0.648 $\pm 0.074$ 0.060 $\pm 0.012$ 0.838 $\pm 0.016$
3 0.768 $\pm 0.022$ 0.669 $\pm 0.033$ 0.062 $\pm 0.005$ 0.853 $\pm 0.013$
$\frac{1}{\sqrt{d}}$ 0 0.604 $\pm 0.134$ 0.464 $\pm 0.115$ 0.046 $\pm 0.005$ 0.735 $\pm 0.066$
1 0.734 $\pm 0.048$ 0.702 $\pm 0.031$ 0.058 $\pm 0.014$ 0.857 $\pm 0.010$
2 0.724 $\pm 0.064$ 0.688 $\pm 0.080$ 0.060 $\pm 0.007$ 0.864 $\pm 0.018$
3 0.762 $\pm 0.017$ 0.690 $\pm 0.062$ 0.057 $\pm 0.012$ 0.850 $\pm 0.022$
Appendix E Full main-setting results
Table 13 reports the main-setting results for the best hyperparameter combinations for all three tasks of predicting tags, edges, and relations of the semantic dependency graphs. The first part of Table 13 is identical to the bottom of Table 3. Since we already discussed the performance for labeled edges in Section 5, we forgo discussing them again in this appendix.
As regards unlabeled edges, using score normalization improves mean performance for all datasets, similarly to relations. This is expected, as unlabeled edges directly influence the prediction of the edge labels, together with entity class prediction. As we have already observed for relations, the performance boost provided by normalization is particularly great for SciERC at $N=0$.
It is also interesting to notice that for CoNLL04, the performance for the unlabeled edge prediction task is only slightly higher than that of the relations. In fact, as regards ADE, the opposite is true, with the labeled performance seemingly being higher. Due to the high variances, it would therefore seem that for these two datasets there is essentially no difference in difficulty between the labeled and unlabeled edge prediction tasks. In the case of ADE, this makes perfect sense, since there is only one possible relation. For CoNLL04, the amount of relation types is rather limited as well; therefore, it is sensible for the performance to also be similar in this case. Indeed, when looking at SciERC and ERFGC, the performance gap between labeled and unlabeled tasks is considerable. This is a strong indication that CoNLL04 is thus found somewhere in the middle, where the number of possible relations only slightly affects labeled performance.
As regards enEWT and SciDTB, the performance on the prediction of edges (UAS) and relations (LAS) is also rather similar. Since in this case the predictions involve syntactic rather than semantic relations, the similar performance hints at relations being easier to predict in SynDP than in SemDP.
In the tagging task, very high standard deviations can be observed. In the case of ADE and CoNLL04, this could be caused by our choice of $\lambda_{1}=0.1$ being too low a value. Regarding SciERC, as already mentioned, most of the entities in the validation and test sets are not found in the training set. This means it is not possible to train the model to recognize them, which results in the abysmal tagging performance reported in Table 13. This makes it challenging to train good tag embeddings which can help inform the edge and relation prediction tasks. Surprisingly, we notice that for CoNLL04 and ERFGC the tagging performance is also seemingly higher with score normalization. However, the overlaps of the standard deviations are too high to be able to make any claims on the matter.
Appendix F Computational resources
We ran all of our experiments on a cluster of NVIDIA H100 (96GB of VRAM) and NVIDIA L40 (48GB of VRAM) GPUs, one run per single GPU. When freezing the BERTbase encoder and only training the BiLSTMs and the classifiers, each training and evaluation run took $\scriptstyle\mathtt{\sim}$5-7 minutes, depending on the number of BiLSTM layers. For the main setting, finding the best hyperparameters involved training and testing 6,120 models, for a total of $\scriptstyle\mathtt{\sim}$600 GPU hours. In the full fine-tuning setting (Appendix D.6), training and evaluation took $\scriptstyle\mathtt{\sim}$1 hour for each of the 40 base models and $\scriptstyle\mathtt{\sim}$2-3 hours for each of the 40 large models, for an additional $\scriptstyle\mathtt{\sim}$140 GPU hours.
@@ 3. Answer: [Yes]
  4. Justification: Our abstract and introduction faithfully report the claims made in the rest of the paper, i.e. we show theoretically and empirically that normalizing biaffine scores for dependency parsing in NLP provides higher parsing performance.
@@ 23. Answer: [Yes]
  24. Justification: The code is provided. The repository includes a README.md file which guides the reader through its usage. The data for ADE, CoNLL04, and SciERC can be downloaded from the link provided in Section 4.1. As regards ERFGC, [46] provide access to the corpus upon request. UD 2.2 enEWT and SciDTB are openly available.
@@ 28. Answer: [Yes]
  29. Justification: Yes, we report information about the data we use in Section 4.1. Training and evaluation information is provided in Sections˜4.2, 4.3 and 4.4.
@@ 35. Guidelines:
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
@@ 38. Answer: [Yes]
  39. Justification: Appendix F provides all relevant information with relation to the hardware used, its characteristics, and the time necessary to reproduce the experiments.
@@ 43. Answer: [Yes]
  44. Justification: Our work does not raise any ethics concerns with regard to any of the guidelines listed within the NeurIPS Code of Ethics. We have no human participants, we use publicly available data, and we believe our work does not have any broader potentially unintended societal impact.
@@ 60. Guidelines:
  63. Answer: [Yes]
  64. Justification: We provide a well-documented GitHub repo, which is linked to in the abstract.